The closed economy model with government is considered. We observe following values of autonomous expenditures: Co = 150, Io = 50, Go = 100, To= 50. Marginal propensity to consume equals 0.8, investment sensitivity parameter equals 5 and income tax rate equals 20%. Using this information find IS curve and derive it graphically.
Answer #1

Pago no Co=150..Io -50, Go = 100, To - 50 MPC = 0.8 Anv. Sensitivity = 5, Income tax rate (t)= 20). Consumption_Function : cDate / – Page no. 2 1 0 .30 0.36 z > Ya entre 200-51] Y-- 100 (260-51] 1 Y= 25 f260-5i | - a Egiuation of . Is curve. * GrapPage ndes Date / 16 If I= 16% then I=49.2 and 4 = 25 (260-5. y= 25 (260- 0.8 y= 258 259.2 · 259.2 Y=720 Summary: When i r